[figure: 9.7 cm]
Ce_{3}Pt_{3}Bi_{4}, Ce_{3}Pt_{3}Sb_{4}, Ce_{3}Au_{3}Sb_{4} AND Ce_{3}Cu_{3}Sb_{4}:

KONDO INSULATORS, KONDO LATTICES, FERROMAGNETS,

METALS OR SEMICONDUCTORS?

P. Wachter, L. Degiorgi, G. Wetzel and H. Schwer Laboratorium für Festkörperphysik, ETH Zürich, 8093 Zürich, Switzerland

Ce_{3}Pt_{3}Bi_{4} was the material which first has been claimed to be a Kondo insulator. There is evidence, especially from the Knight shift that the magnetic susceptibility is zero for zero temperature, but there is a peak near 100 K. That is, Ce must be in a nonmagnetic state for T → 0, namely Ce^{4+}. As well from resistivity as from optical reflectivity a gap of several meV has been deduced. The Luttinger theorem permits only the Fermi level in a gap with an even f − d count, which is only the case for Ce^{4+}. The proximity of an empty 4f^{1} state above, but close to E_{F} leads to f − d hybridization. Pt is electrically neutral in this compound. Similar, but not so pronounced in its effects is Ce_{3}Pt_{3}Sb_{4}. In Ce_{3}Au_{3}Sb_{4} Au donates one electron to Sb, so Ce is Ce^{3+}, the material is antiferromagnetic and has no gap, at most a zero gap. In Ce_{3}Cu_{3}Sb_{4} Cu donates also one electron to Sb, so Ce is again trivalent, and the material was claimed to be ferromagnetic, but it only has a spontaneous moment of 0.66 μ_{B}/Ce and is probably a canted antiferromagnet. It was also claimed to be a semiconductor, but resistivity, Hall effect and optical reflectivity clearly show that no gap exists.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : PLENARY LECTURE

Corresponding author : Prof. Dr. P. Wachter

Address for correspondence : Laboratorium für Festkörperphysik ETH Zürich 8093 Zürich Switzerland

Email address : wachter@solid.phys.ethz.ch